Find the slope of the line that contains (-1,9) and (-1, 1)
Answer by sudhanshu_kmr(1152)


slope of line = 9-1 / -1 +1 = 8/0 = not defined...

so, line is parallel to y axis ,
slope = tan90 =not defined